An emf is induced in a conducting loop of wire 1.06 m long as its shape is changed from square to circular.

Part A

Find the average magnitude of the induced emf if the change in shape occurs in 4.06 s and the local 0.117-Tmagnetic field is perpendicular to the plane of the loop.
Answer #1

?T = 4.06 s
B = 0.117 T
L = 1.06 m

A_square = (L / 4) ^ 2
= 0.070 m^2

c = 1.06 m
r = 1.06 / (2 * pi)
A_circle = pi * r ^ 2
=0.089 m^2

?A = A_square - A_circle =-0.019

EMF = - d? /dt
EMF = - B (?A / ?T)
EMF = 5.47*10-4 V